OpenStudy (anonymous):

anyone help.... i'm stuck at doing cotx= 3/2. how you do it? I replaced cotx with cosx/sinx, but i end up not knowing how to solve it. please help! :D

OpenStudy (anonymous):

try writing \[cotx = \frac{1}{tanx}\]

OpenStudy (anonymous):

\[\frac{1}{tanx} = \frac{3}{2} => 1 = \frac{3}{2}tanx => \frac{3}{2}tanx = 1 => 3 tanx = 2 => tanx = \frac{2}{3}\]

OpenStudy (anonymous):

now solve for \(x\)

OpenStudy (anonymous):

got it. 33.7 :) thanks!